OpenStudy (anonymous):

Purvi and her friends tipped their waiter a total of 6for dinners that cost $8,$9, $11, $12 and each. Use a percent proportion to solve for the percent tip they left the waiter. Show your work.

OpenStudy (anonymous):

6/(8+9+11+12)=6/40=15%